# approximate nearest neighbour IOU intersection over union

### v1

@@ 4.2 Baselines
Additionally, we compare against the multiview-consistent, but not material-aware, SAM2 model. Finally, “Ours” denotes our full method, including our fine-tuned network. For all methods, we lift the results to 3D using our point-cloud representation.
@@ 4.3 Results
We evaluate each method in 3D (after the point cloud lookup), along three axes: selection quality, robustness and multiview consistency. We report metrics on binary selection masks obtained by thresholding similarities against 0.5. Metrics are normalized to $[0,1]$; see Appendix B for details.
@@
Table 1: Selection accuracy across datasets (columns) for several methods (rows), with 95% confidence intervals. For the per-scene measurements and precision and recall, we refer to Appendix B. Mat. is short for Materialistic [56].
@@
Table 2: Multiview consistency (top) and robustness (bottom) of our selection across unseen test views. We report Hamming distance (×$\times 100$) with 95% confidence intervals. Lower is better.
Figure 7: Selection results on NeRFs, 3D Gaussians and meshes across objects (columns) and methods (rows). Leftmost column shows the clicked view and the click. Selection results, from a novel view, are overlaid as green masks.
@@
We perform a click in one view, then for each of 50 random novel views we compare the obtained selection mask against a rendered ground-truth mask. We compute mean intersection over union (mIoU), a classical selection metric measuring the overlap between the masks. We also report F1 score which is the harmonic mean of precision and recall, and is more robust than either alone. We average each metric over the views and over five random clicks on each material. We report the averages and 95% confidence intervals across the datasets in Tab. 1, higher is better. Appendix B provides a per-scene breakdown.
@@
We show selection results on all our evaluated modalities (NeRFs, 3D Gaussians, meshes) in Fig. 7. We see that Materialistic and Materialistic MV do not work well on high-frequency boundaries of objects and that SAM2 cannot be trivially used for material selection, with parts of or entire objects selected at once. In contrast, our method creates sharp boundaries, including around thin elements, selecting parts of the object with materials similar to that clicked in the first view, and is robust to lighting variations (see Fig. 8).
We further evaluate our results in 2D without the point cloud lookup, which improves average mIoU by around 5%. However, our 3D aggregation in a point cloud lookup provides a significant advantage in efficiency, reducing per-frame inference processing times from around 5 sec in 2D to around 10 ms in 3D (500$\times$ faster), making it a more practical choice overall. This difference in quality is mainly explained by the depth estimation quality on volumetric representations, which is not always perfect.
@@ 5.1 Segmentation
While our method targets selection, it can be used to automatically segment an object into material subparts. Inspired by the image-level sampling in SAM [31], we propose an equivalent approach for materials on 3D objects. It involves two steps: (1) automatically choosing “selection clicks” and (2) merging similar resulting selections.
Densely sampling an entire object from multiple views is impractical (500-click sampling of the Lego asset in Fig. 3 takes $\sim$20 min). Instead, we select clicks based on the modes of a CIE LAB (D65) 3D histogram. To account for color differences due to shading and illumination, we discretize the L channel more coarsely than the AB channels (4 vs 16 bins). This strategy places samples in locations of different color, which often coincide with material regions. We sample 25 points in total, with stratified sampling in each color mode area, proportionally to the color area.
Given the sampled clicks, we compute the material similarities as outlined in Sec. 3 for a random set of views, to obtain a binary selection masks per click. We compare these masks to one another and store their pairwise mIoU in a (symmetric) matrix, where each entry represents how similar the selection for two different clicks is. A high value implies that the clicks which created this entry led to similar selections, i.e., were on the same material, so we can safely keep only one of them. We repeat this process until all matrix entries are below the empirically determined threshold 0.75, leaving only clicks on truly different materials. We show results of this segmentation approach in Fig. 9.
@@ 5.2 Editing
NeRFs.
For NeRFs, we demonstrate color editing. We ray-march the NeRF as usual, but for each 3D point we query whether it has been selected. If yes, we adjust the color returned by the NeRF through a color shift in LAB space, to preserve relative shading and lighting information. We show an example in LABEL:fig:Teaser and in the appendix.
Material-aware 3D Gaussians.
For 3D Gaussians, we use our material segmentation step (described in Sec. 5.1). We then render the respective material masks for each training view and convert them to ID masks, so each pixel in the training images is associated with a material index. We then re-train the Gaussians with an extra channel for materials which is treated like the RGB channels for rasterization. This creates a clear separation between Gaussians at material boundaries, simplifying downstream edits, and provides a per-Gaussian material handle. We can now select all Gaussians that encode, e.g., material number two, and edit their properties such as color, position or density. In LABEL:fig:Teaser we move the gazebo’s wooden base upwards and set the white painted regions’ density to zero.
Meshes.
For selection on meshes, we exploit their UV parametrization by writing the selected material similarities to a 2D UV map. This enables trivial creation of material-ID maps, or changing the selected material. Here, because the similarities are directly projected to pixel values, we find it beneficial to use the hole-filling and sprinkle-removal techniques described in the original SAM2 paper [49].
We show results on the output of text-to-mesh generated assets and other meshes [65]. Using our automatic segmentation we can easily replace the diffuse textures on a text-to-3D generated asset with PBR materials.
Figure 9: Automatic segmentation of meshes into materials.
@@
We find our method to significantly improve material selection in 3D, however some limitations remain to be addressed. Selecting materials on objects like glass and mirror remains a challenge as it is unclear if a user would prefer to select the transparent/mirror material or what is behind/reflected. Our method also depends on precise 3D reconstruction for accurate material selection. Errors in depth reconstruction can cause noise in our point clouds and inaccurate lookups in novel views. Improving depth estimation in volumetric reconstruction will help mitigate this issue.
@@ Conclusion
We present SAMa, a material selection model for 3D, leveraging a video model for cross-view consistency and a simple yet efficient projection to 3D in the form of a similarity point cloud. Our approach enables interactive material selection, visualization and downstream manipulation of the 3D assets. As we specialize the SAM2 video model to a new modality, we find that finetuning using videos is important and that 500 varied videos are enough to change the modality. We believe this opens interesting opportunities to explore selection across various modalities.

### v2

@@ 4.2 Baselines
Additionally, we compare against the multiview-consistent, but not material-aware, SAM2. Finally, “Ours” denotes our full method, including our fine-tuned network. For all methods, we lift the results to 3D using our point-cloud representation. We show additional comparisons [34, 27, 42, 9] in Suppl. Figs. 11 – 14.
@@ 4.3 Results
We evaluate each method in 3D (after the point cloud lookup), along three axes: selection quality, robustness and multiview consistency. We report metrics on binary selection masks obtained by thresholding similarities against 0.5. Metrics are normed to $[0,1]$; see Suppl. Sec. B for details.
@@
Table 1: Selection accuracy across datasets (columns) for several methods (rows), with 95% confidence intervals. For the per-scene measurements and precision and recall, we refer to Suppl. Sec. B. Mat. is short for Materialistic [63].
@@
Table 2: Multiview consistency (top) and robustness (bottom) of our selection across unseen test views. We report Hamming distance ($\times 100$) with 95% confidence intervals. Lower is better.
Figure 7: Selection results on NeRFs, 3D Gaussians and meshes across objects (columns) and methods (rows). The top row shows the clicked view and the click, while selection results, from a novel view, are overlaid as green masks in the lower rows. We show results on assets from standard NeRF/3DGS benchmarks as well as a generated [72] mesh (cup) and photogrammetry capture (birdhouse), and more in supplemental material.
@@
We perform a click in one view, then for each of 50 random novel views we compare the obtained selection mask against a rendered ground-truth mask. We compute mean intersection over union (mIoU), a classical selection metric measuring the overlap between the masks. We also report F1 score which is the harmonic mean of precision and recall, and is more robust than either alone. We average each metric over the views and over five random clicks on each material. We report the averages and 95% confidence intervals across the datasets in Tab. 1, higher is better. Suppl. Sec. B provides a per-scene breakdown.
@@
We show selection results on all our evaluated modalities (NeRFs, 3D Gaussians, meshes) in Fig. 7. We see that Materialistic and Materialistic MV do not work well on high-frequency boundaries of objects and that SAM2 cannot be trivially used for material selection, with only parts or entire objects selected at once. In contrast, our method creates sharp boundaries, including around thin elements, and is robust to lighting variations (see Fig. 8). For additional results on generated- as well as real-world meshes and neural assets, please see the supplemental material. We also include comparisons to MatSeg3D [42], which focuses on mesh segmentation into semantic material classes, and methods focusing on object segmentation and selection in 3DGS [27, 77, 34].
[…]
@@ 5.1 Segmentation
While our method targets selection, we propose an automatic segmentation mode to divide an object into material subparts inspired by the image-level sampling in SAM [35].
Densely sampling an entire object from multiple views is impractical (500-click sampling of the Lego asset in Fig. 2 takes $\sim$20 min). Instead, we propose a light-weight, greedy alternative. We first randomly sample a click on the asset, for which we compute the material similarity (as outlined in Sec. 3) in a fixed set of 15 views (elevation in $\left[-30\degree,0\degree,30\degree\right]$, azimuth in five 72°  increments). We binarize these per-view selections and project them to UV space, assigning a unique ID to the clicked material. We then sample a new point in an area that has not yet been segmented.
We repeat this process until the entire mesh surface is segmented, which yields an ID map in UV space in which each point of the visible mesh surface corresponds to a unique material ID via its UV coordinate. We can use this map to edit specific material regions, for PBR replacement, or to segment the mesh along the material ID regions. With this approach, the segmentation of a complete mesh takes under 10 seconds. We show results of our method in Fig. 9.
@@ 5.2 Editing
NeRFs. For NeRFs, we demonstrate color editing. We ray-march the NeRF as usual, but for each 3D point we query whether it has been selected. If yes, we adjust the color returned by the NeRF through a color shift in LAB space, to preserve relative shading and lighting information. We show an example in LABEL:fig:Teaser and in Suppl. Fig. 7.
Material-aware 3D Gaussians. For 3D Gaussians, we use our material segmentation step (described in Sec. 5.1). We then render the respective material masks for each training view and convert them to ID masks, so each pixel in the training images is associated with a material index. We then re-train the Gaussians with an extra channel for materials which is treated like the RGB channels for rasterization. This creates a clear separation between Gaussians at material boundaries, simplifying downstream edits, and provides a per-Gaussian material handle. We can now select all Gaussians that encode, e.g., material number two, and edit their properties such as color, position or density. In LABEL:fig:Teaser we move the gazebo’s wooden base upwards and set the white painted regions’ density to zero.
Meshes. For selection on meshes, we exploit their UV parametrization by writing the selected material similarities to a 2D UV map. This enables trivial creation of material-ID maps, or change of a selected material. Here, because the similarities are directly projected to pixel values, we find it beneficial to use the hole-filling and sprinkle-removal techniques described in the original SAM2 paper [56].
We show results on the output of text-to-mesh generated assets [72]. Using our automatic segmentation we can easily replace the diffuse textures on a text-to-3D generated asset with PBR materials (see Suppl. Fig. 10).
Figure 9: Automatic segmentation of synthetic ($3^{rd}$) and generated [72] ($1^{st}$, $2^{nd}$, $4^{th}$) meshes into material ID maps.
@@
We find our method to significantly improve material selection in 3D. However, some limitations remain to be addressed; for instance, selecting materials on objects like glass and mirrors remains challenging as it is unclear if a user would prefer to select the transparent/mirror material or what is behind/reflected. Similarly, our method currently assumes a fixed definition of materials and cannot adapt to user-defined preferences, such as focusing solely on specific parameters like roughness.
Further, SAMa depends on precise 3D reconstruction for accurate material selection. Errors in depth reconstruction can cause noise in our point clouds and inaccurate lookups in novel views, an issue which can be mitigated by improving depth estimation in volumetric reconstruction [74]. Finally, unseen parts of meshes are selected based on their nearest visible 3D points, which could be improved with a UV space propagation of the selection [12].
@@ Conclusion
We present SAMa, a material selection model for 3D, leveraging a video model for cross-view consistency and a simple yet efficient projection to 3D in the form of a similarity point cloud. Our approach enables interactive material selection, visualization and downstream manipulation of the 3D assets at high quality. As we specialize the SAM2 video model to a new modality, we find that video-based finetuning is crucial, and that 500 varied videos are sufficient to change the modality. We believe this opens interesting opportunities to explore selection across various modalities.